# Distributed Retraction-Free and Communication-Efficient Optimization on the Stiefel Manifold

### v1

@@ 1 Introduction
$\min_{X\in\mathbb{R}^{n\times p}}$ $f(X)=\frac{1}{N}\sum_{i=1}^{N}\big{[}f_{i}(X):=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}F(X;\xi_{i})\big{]},$ (1a)
$\mathrm{s.\,t.}\quad$ $X^{\top}X=I_{p},$ (1b)
@@ 3.2 Compressor and Error Feedback
Suppose $\nabla f(X^{k})\to\nabla f(X^{\star})$ as $k\to\infty$, where $X^{\star}$ is a stationary solution to problem (4). The above inequality implies that $Y^{k}$ converges to $\nabla f(X^{\star})$. Combined with (5a), this guarantees that $X^{k}\to X^{\star}$ under compression, thereby validating the effectiveness of the error feedback method.
@@ 4.3 EF-Landing Algorithm
Scenarios Deterministic Stochastic (with momentum)
With compression EF-Landing (Det.$\sharp$) $\sim\mathcal{O}(1/K)$ EF-Landing (Sto.$\flat$) $\sim\mathcal{O}(1/\sqrt{NK})$
Without compression Vanilla Landing $\sim\mathcal{O}(1/K)$ Stochastic Landing $\sim\mathcal{O}(1/\sqrt{NK})$
@@ 5.1 Assumptions
$\|\nabla F(X;\xi_{i})\|_{F}\leq L^{\prime},\quad\forall X\in\mathrm{St}(p,n)^{\epsilon},\,\forall\xi\sim\mathcal{D}_{i}.$
@@ 5.2 Supporting Lemmas
[…]
@@
[…]
@@ 5.3 Main Convergence Theorem
$\mathcal{L}^{k}:=m(X^{k})-m^{\star}+\frac{c_{1}\gamma}{\theta}\tilde{G}^{k}+\frac{2c_{1}\gamma\eta\beta}{\theta}\tilde{P}^{k}+\frac{c_{2}\gamma}{\eta}P^{k}$ (12)
@@
[…]
@@
  • When the momentum coefficient $\eta=1$ and the variance of gradients $\sigma^{2}=0$, the momentum and stochastic error is 00, and $c_{1},c_{2}$ are specified as $c_{1}=1,c_{2}=0$.
  • When the contractive factor $\alpha=1$, the compression error is 00, and $c_{1},c_{2}$ are specified as $c_{1}=0,c_{2}=1$.
@@ 5.4 Convergence Rate in Deterministic Scenario
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\mathrm{grad}f(X^{k})\|_{F}^{2}]\leq~{}$ $\frac{4(m(X^{0})-m^{\star})}{K\gamma}+\frac{4\mathbb{E}[\tilde{G}^{0}]}{K\theta},$
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\mathcal{N}(X^{k})]\leq~{}$ $\frac{2(m(X^{0})-m^{\star})}{K\gamma\lambda\mu}+\frac{2\mathbb{E}[\tilde{G}^{0}]}{K\theta\lambda\mu},$
@@
$\frac{1}{K}\sum_{k=1}^{K-1}\mathcal{N}(X^{k})\leq~{}$ $\frac{2(m(X^{0})-m^{*})}{K\gamma\lambda\mu}.$
@@ 5.5 Convergence Rate in Stochastic Scenario
$\frac{1}{K}\sum_{k=1}^{K-1}\mathbb{E}[\|\mathrm{grad}f(X^{k})\|_{F}^{2}]\leq~{}$ $4\mathcal{O}\Big{(}\frac{1}{\sqrt{NK}}\Big{)},$
$\frac{1}{K}\sum_{k=1}^{K-1}\mathbb{E}[\mathcal{N}(X^{k})]\leq~{}$ $\frac{2}{\lambda\mu}\mathcal{O}\Big{(}\frac{1}{\sqrt{NK}}\Big{)},$
@@
$\frac{1}{K}\sum_{k=1}^{K-1}\mathcal{N}(X^{k})\leq~{}$ $\frac{2C^{\gamma}_{1}\mathcal{L}^{0}}{K\lambda\mu}+\frac{2\sigma}{\lambda\mu}\sqrt{\frac{C^{\gamma}_{2}\mathcal{L}^{0}}{NK}},$
@@ Appendix A Details of Realizing EF-Landing
[…]
@@ Appendix B Details of Convergence Analysis
[…]
@@
[…]
@@
[…]
@@
  • When the momentum rate $\eta=1$ and the variance of gradients $\sigma^{2}=0$, the momentum and stochastic error will be 00, so $c_{1},c_{2}$ are specified as $c_{1}=1,c_{2}=0$, which is trivial.
  • When the contractive factor $\alpha=1$, the compression error will be 00, so $c_{1},c_{2}$ are specified as $c_{1}=0,c_{2}=1$, which is trivial too.
@@
STEP 4. Summing Up. Summing up all inequalities from $k=0$ to $K-1$ and dividing by $K$, we have
@@
2. Theorem 5.12. When $\alpha=1$, the compression error $\tilde{G}^{k}$ will be 00, so the Lyapunov function can be further simplified as $\mathcal{L}^{k}=m(X^{k})-m^{*}$. And the choice of the step size can also be simplified as
@@ B.5 An adequate Selection of Step Size with momentum
$\frac{1}{4K}\sum_{k=1}^{K-1}\mathbb{E}[\|\mathrm{grad}f(X^{k})\|_{F}^{2}]+\frac{\lambda\mu}{2K}\sum_{k=1}^{K-1}\mathbb{E}[\mathcal{N}(X^{k})]\leq\frac{C^{\gamma}_{1}\mathcal{L}^{0}}{K}+\frac{C^{\gamma}_{2}\mathcal{L}^{0}}{\eta K}+\frac{2\eta^{2}(1-\alpha)\sigma^{2}}{\theta}+\frac{4\eta^{3}\beta\sigma^{2}}{\theta}+\frac{2\eta\sigma^{2}}{N}.$ (68)
@@
$\frac{1}{4K}\sum_{k=1}^{K-1}\mathbb{E}[\|\mathrm{grad}f(X^{k})\|_{F}^{2}]+\frac{\lambda\mu}{2K}\sum_{k=1}^{K-1}\mathbb{E}[\mathcal{N}(X^{k})]\leq\frac{C^{\gamma}_{1}\mathcal{L}^{0}}{K}+4\frac{C^{\gamma}_{2}\mathcal{L}^{0}}{\eta K}$
@@
$\frac{1}{4K}\sum_{k=1}^{K-1}\mathbb{E}[\|\mathrm{grad}f(X^{k})\|_{F}^{2}]+\frac{\lambda\mu}{2K}\sum_{k=1}^{K-1}\mathbb{E}[\mathcal{N}(X^{k})]\leq\frac{\mathcal{L}^{0}}{\gamma K}+\frac{\eta\sigma^{2}}{N}.$ (71)
@@
$\frac{1}{4K}\sum_{k=1}^{K-1}\mathbb{E}[\|\mathrm{grad}f(X^{k})\|_{F}^{2}]+\frac{\lambda\mu}{2K}\sum_{k=1}^{K-1}\mathbb{E}[\mathcal{N}(X^{k})]\leq\frac{C^{\gamma}_{1}\mathcal{L}^{0}}{K}+\frac{C^{\gamma}_{2}\mathcal{L}^{0}}{\eta K}+\frac{\eta\sigma^{2}}{N}.$ (73)
@@ C.2 Convergence Results of EF-Landing in Stochastic Scenarios for Block-wise Constraints
$\frac{1}{K}\sum_{k=1}^{K-1}\sum_{j=1}^{J}\mathbb{E}[\|\mathrm{grad}_{j}f(X_{j}^{k})\|_{F}^{2}]\leq$ $~{}4\mathcal{O}\Big{(}\frac{1}{\sqrt{NK}}\Big{)},$
$\frac{1}{K}\sum_{k=1}^{K-1}\sum_{j=1}^{J}\mathbb{E}[\mathcal{N}_{j}(X_{j}^{k})]\leq$ $~{}\frac{2}{\lambda\mu}\mathcal{O}\Big{(}\frac{1}{\sqrt{NK}}\Big{)},$
$\frac{1}{K}\sum_{k=1}^{K-1}\mathbb{E}\Big{[}\Big{\|}\frac{\partial f}{\partial x^{k}}\Big{\|}_{2}^{2}\Big{]}\leq$ $~{}2\mathcal{O}\Big{(}\frac{1}{\sqrt{NK}}\Big{)},$
@@
Summing up all inequalities from $k=0$ to $K-1$ and dividing by $K$, we have

### v2

@@ 1 Introduction
$\min_{X\in\mathbb{R}^{n\times p}}\$$f(X)\hskip-1.42262pt=\hskip-1.42262pt\frac{1}{N}\hskip-1.42262pt\sum_{i=1}^{N}\hskip-1.42262pt\big{[}f_{i}(X)\hskip-1.42262pt:=\hskip-1.42262pt\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}F(X;\xi_{i})\big{]},$ (1a)
$\mathrm{s.\,t.}\quad\$$X^{\top}X\hskip-1.42262pt=\hskip-1.42262ptI_{p},$ (1b)
@@ 3.2 Compressor and Error Feedback
Suppose $\nabla f(X^{k})\to\nabla f(X^{*})$ as $k\to\infty$, where $X^{*}$ is a stationary solution to problem (4). The above inequality implies that $Y^{k}$ converges to $\nabla f(X^{*})$. Combined with (5a), this guarantees that $X^{k}\to X^{*}$ under compression, thereby validating the effectiveness of the error feedback method.
@@ 4.3 EF-Landing Algorithm
Scenarios
Deterministic
Stochastic
(with momentum)
With
compression
EF-Landing
[…]
$\sim\mathcal{O}(1/K)$
EF-Landing
[…]
$\sim\mathcal{O}(1/\sqrt{NK})$
Without
compression
Vanilla
Landing
$\sim\mathcal{O}(1/K)$
Stochastic
Landing
$\sim\mathcal{O}(1/\sqrt{NK})$
@@ 5.1 Assumptions
$\|\nabla F(X;\xi_{i})\|_{F}\leq L^{\prime},\quad\forall X\in\mathrm{St}(p,n)^{\epsilon},\,\forall\xi_{i}\sim\mathcal{D}_{i}.$
@@ 5.2 Supporting Lemmas
$\gamma\leq\gamma_{s}=\min\Big{\{}$ $\frac{\lambda(1-\epsilon)\epsilon}{(1+\epsilon)^{2}(L^{\prime})^{2}+\lambda^{2}(1+\epsilon)\epsilon^{2}},$
@@
$+$~ $\frac{1}{4}\|\mathrm{skew}(\boldsymbol{g}X^{\top})X\|_{F}^{2}-\|\boldsymbol{g}-\nabla f(X)\|_{F}^{2},$
@@ 5.3 Main Convergence Theorem
$\mathcal{L}^{k}\hskip-2.27621pt:=\hskip-2.27621ptm(X^{k})\hskip-1.42262pt-\hskip-2.27621ptm^{*}\hskip-2.27621pt+\hskip-2.27621pt\frac{c_{1}\gamma}{\theta}\tilde{G}^{k}\hskip-2.27621pt+\hskip-2.27621pt\frac{2c_{1}\gamma\eta\beta}{\theta}\tilde{P}^{k}\hskip-2.27621pt+\hskip-2.27621pt\frac{c_{2}\gamma}{\eta}P^{k}$ (12)
@@
$\leq$~ $\frac{\mathcal{L}^{0}}{\gamma K}+\frac{c_{1}\eta^{2}(1-\alpha)\sigma^{2}}{\theta}+\frac{2c_{1}\eta^{3}\beta\sigma^{2}}{\theta}+\frac{c_{2}\eta\sigma^{2}}{N},$
@@
  • When the momentum coefficient $\eta=1$ and the variance of gradients $\sigma^{2}=0$, the momentum and stochastic error is 0, and $c_{1},c_{2}$ are specified as $c_{1}=1,c_{2}=0$.
  • When the contractive factor $\alpha=1$, the compression error is 0, and $c_{1},c_{2}$ are specified as $c_{1}=0,c_{2}=1$.
@@ 5.4 Convergence Rate in Deterministic Scenario
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\mathrm{grad}f(X^{k})\|_{F}^{2}]\leq$~ $\frac{4(m(X^{0})-m^{*})}{K\gamma}+\frac{4\mathbb{E}[\tilde{G}^{0}]}{K\theta},$
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\mathcal{N}(X^{k})]\leq$~ $\frac{2(m(X^{0})-m^{*})}{K\gamma\lambda\mu}+\frac{2\mathbb{E}[\tilde{G}^{0}]}{K\theta\lambda\mu},$
@@
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{N}(X^{k})\leq$~ $\frac{2(m(X^{0})-m^{*})}{K\gamma\lambda\mu}.$
@@ 5.5 Convergence Rate in Stochastic Scenario
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\mathrm{grad}f(X^{k})\|_{F}^{2}]\leq$~ $4\mathcal{O}\Big{(}\frac{1}{\sqrt{NK}}\Big{)},$
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\mathcal{N}(X^{k})]\leq$~ $\frac{2}{\lambda\mu}\mathcal{O}\Big{(}\frac{1}{\sqrt{NK}}\Big{)},$
@@
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{N}(X^{k})\leq$~ $\frac{2C^{\gamma}_{1}\mathcal{L}^{0}}{K\lambda\mu}+\frac{2\sigma}{\lambda\mu}\sqrt{\frac{C^{\gamma}_{2}\mathcal{L}^{0}}{NK}},$
@@ Appendix A Details of Realizing EF-Landing
$=$~ $\mathrm{Tr}\big{(}\mathrm{skew}(\boldsymbol{g}X^{\top})^{\top}\cdot\lambda X(X^{\top}X-I_{p})X^{\top}\big{)}=\langle\mathrm{skew}(\boldsymbol{g}X^{\top}),\lambda X(X^{\top}X-I_{p})X^{\top}\rangle,$
@@ Appendix B Details of Convergence Analysis
$=$~ $\langle\mathrm{skew}(\boldsymbol{g}X^{\top}),\nabla f(X)X^{\top}\rangle+\lambda\langle X^{\top}X-I_{p},X^{\top}\nabla f(X)\rangle$
$=$~ $\langle\mathrm{skew}(\boldsymbol{g}X^{\top}),\mathrm{skew}(\nabla f(X)X^{\top})\rangle+\lambda\langle X^{\top}X-I_{p},\mathrm{sym}(X^{\top}\nabla f(X))\rangle.$ (19)
@@
$=$~ $\frac{1}{2}\|\mathrm{skew}(\boldsymbol{g}X^{\top})\|_{F}^{2}+\frac{1}{2}\|\mathrm{skew}(\nabla f(X)X^{\top})\|_{F}^{2}-\frac{1}{2}\|\mathrm{skew}((\boldsymbol{g}-\nabla f(X))X^{\top})\|_{F}^{2}$
$\geq$~ $\frac{1}{2}\|\mathrm{skew}(\boldsymbol{g}X^{\top})\|_{F}^{2}+\frac{1}{2}\|\mathrm{skew}(\nabla f(X)X^{\top})\|_{F}^{2}-\frac{\sigma_{1}^{2}}{2}\|\boldsymbol{g}-\nabla f(X)\|_{F}^{2},$ (20)
@@
$=$ $~\mathbb{E}\bigg{[}\mathbb{E}_{\xi_{i}^{k+1}}\Big{[}\big{\|}(1-\eta)\big{(}\boldsymbol{v}_{i}^{k}-\nabla f_{i}(X^{k+1})\big{)}+\eta\big{(}\nabla F(X^{k+1};\xi_{i}^{k+1})-\nabla f_{i}(X^{k+1})\big{)}\big{\|}_{F}^{2}\Big{|}\mathcal{F}^{k+1}_{P}\Big{]}\bigg{]}$
@@
  • When the momentum rate $\eta=1$ and the variance of gradients $\sigma^{2}=0$, the momentum and stochastic error will be 0, so $c_{1},c_{2}$ are specified as $c_{1}=1,c_{2}=0$, which is trivial.
  • When the contractive factor $\alpha=1$, the compression error will be 0, so $c_{1},c_{2}$ are specified as $c_{1}=0,c_{2}=1$, which is trivial too.
@@
STEP 4. Summing Up. Summing up all inequalities from $k=0$ to K1K-1 and dividing by $K$, we have
@@
2. Theorem 5.12. When $\alpha=1$, the compression error $\tilde{G}^{k}$ will be 0, so the Lyapunov function can be further simplified as $\mathcal{L}^{k}=m(X^{k})-m^{*}$. And the choice of the step size can also be simplified as
@@ B.5 An adequate Selection of Step Size with momentum
$\frac{1}{4K}\sum_{k=0}^{K-1}\mathbb{E}[\|\mathrm{grad}f(X^{k})\|_{F}^{2}]+\frac{\lambda\mu}{2K}\sum_{k=0}^{K-1}\mathbb{E}[\mathcal{N}(X^{k})]\leq\frac{C^{\gamma}_{1}\mathcal{L}^{0}}{K}+\frac{C^{\gamma}_{2}\mathcal{L}^{0}}{\eta K}+\frac{2\eta^{2}(1-\alpha)\sigma^{2}}{\theta}+\frac{4\eta^{3}\beta\sigma^{2}}{\theta}+\frac{2\eta\sigma^{2}}{N}.$ (68)
@@
$\frac{1}{4K}\sum_{k=0}^{K-1}\mathbb{E}[\|\mathrm{grad}f(X^{k})\|_{F}^{2}]+\frac{\lambda\mu}{2K}\sum_{k=0}^{K-1}\mathbb{E}[\mathcal{N}(X^{k})]\leq\frac{C^{\gamma}_{1}\mathcal{L}^{0}}{K}+4\frac{C^{\gamma}_{2}\mathcal{L}^{0}}{\eta K}$
@@
$\frac{1}{4K}\sum_{k=0}^{K-1}\mathbb{E}[\|\mathrm{grad}f(X^{k})\|_{F}^{2}]+\frac{\lambda\mu}{2K}\sum_{k=0}^{K-1}\mathbb{E}[\mathcal{N}(X^{k})]\leq\frac{\mathcal{L}^{0}}{\gamma K}+\frac{\eta\sigma^{2}}{N}.$ (71)
@@
$\frac{1}{4K}\sum_{k=0}^{K-1}\mathbb{E}[\|\mathrm{grad}f(X^{k})\|_{F}^{2}]+\frac{\lambda\mu}{2K}\sum_{k=0}^{K-1}\mathbb{E}[\mathcal{N}(X^{k})]\leq\frac{C^{\gamma}_{1}\mathcal{L}^{0}}{K}+\frac{C^{\gamma}_{2}\mathcal{L}^{0}}{\eta K}+\frac{\eta\sigma^{2}}{N}.$ (73)
@@ C.2 Convergence Results of EF-Landing in Stochastic Scenarios for Block-wise Constraints
$\frac{1}{K}\sum_{k=0}^{K-1}\sum_{j=1}^{J}\mathbb{E}[\|\mathrm{grad}_{j}f(X_{j}^{k})\|_{F}^{2}]\leq$ $~4\mathcal{O}\Big{(}\frac{1}{\sqrt{NK}}\Big{)},$
$\frac{1}{K}\sum_{k=0}^{K-1}\sum_{j=1}^{J}\mathbb{E}[\mathcal{N}_{j}(X_{j}^{k})]\leq$ $~\frac{2}{\lambda\mu}\mathcal{O}\Big{(}\frac{1}{\sqrt{NK}}\Big{)},$
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\Big{[}\Big{\|}\frac{\partial f}{\partial x^{k}}\Big{\|}_{2}^{2}\Big{]}\leq$ $~2\mathcal{O}\Big{(}\frac{1}{\sqrt{NK}}\Big{)},$
@@
Summing up all inequalities from $k=0$ to K1K-1 and dividing by $K$, we have